# Extension principles for the Einstein Yang–Mills system

### v1

@@ 1.3. Outline of this paper
This paper is organized as follows. In section 2, we formulate the SSEYM system. Section 3 is devoted to establishing the extension theorem and Cauchy stability in $H^{1}$ space. The counterparts in the weighted $H^{1}$ space are proved in Section 4. We also demonstrate the local existence theorem in Appendix A, the extension principle away from the axis in Appendix B, and provide a density theorem in Appendix C.
@@ 1.4. Acknowledgement
J.L. is supported by National Key R&D Program of China (No. 2022YFA1005400) and NSFC (12326602, 12141106). J.W. is supported by NSFC (Grant No. 12271450).
@@ 2.1. Spherically symmetric Einstein Yang–Mills equations
In this section, we shall formulate the SSEYM system with certain gauge assumptions.
On a spherical symmetric spacetime $(\mathcal{M},g)$, there exists an isometric action of the rotational group $SO(3)$, where the orbits are $2$-dimensional spheres. Define the axis $\Gamma$ of $(\mathcal{M},g)$ to be the set of fixed points of the $SO(3)$ action. Let us introduce the double null coordinates $(u,\underline{u})$, and denote the level sets by $C_{u}$ and $\underline{C}_{\underline{u}}$, which are the incoming and outgoing null cones. The intersection of $C_{u}$ and $\underline{C}_{\underline{u}}$ is a $2$-sphere denoted by $S_{\underline{u},u}$. We define the area function $r$ by
@@
The metric on $\mathcal{M}$ takes the form of
@@
where $\mathrm{d}\sigma_{S^{2}}$ is the standard metric on $S^{2}$ and $\Omega$ is the lapse function such that $-2\Omega^{2}=g(\partial_{u},\partial_{\underline{u}})$.
@@
with $[\tau_{i},\,\tau_{j}]=\epsilon_{ijk}\tau_{k}$, $\epsilon_{ijk}$ being the Levi-Civita symbol, $\{i,j,k\}\subset\{1,2,3\}$. We assume the spherically symmetric purely magnetic ansatz [4, 6] for the Yang–Mills potential,
@@
[…]
As followed from the ansatz (2.2), $F_{u\underline{u}}=0$, namely the electric charge vanishes; $F_{AB}=\frac{w^{2}-1}{r^{2}}\tau_{3}\varepsilon_{AB}$ [^3] The capital Latin letters $\{A,B\}$ refer to indices on $S^{2}$. and we define the magnetic charge within the sphere $S_{\underline{u},u}$ to be
@@
The vacuum of the Yang–Mills field are given by $w=\pm 1$.
The Einstein Yang–Mills system takes the form of
@@
[…]
@@
We are now ready to write down the spherical symmetric Einstein Yang–Mills system (SSEYM). With the ansatz (2.2) and the double gauge, the SSEYM system consistes of the following equations for $(r,\Omega,w)$,
@@
A characteristic initial data set for the SSEYM system composes of a triplet $(r,\Omega,\partial_{r}w)|_{C_{u_{0}}}$ defined on an outgoing cone $C_{u_{0}}$, see Remark 1.1. In addition to the initial condition, the SSEYM system is incorporated with the following boundary condition
@@
Note that, $Q|_{\Gamma}=0$ ensures the regularity of (2.5) on the axis.
We can reformulate the SSEYM system by eliminating $\Omega$ and replacing it by $m$, so that it is equivalent to the system of $(r,m,w)$ which is composed of
@@
with the boundary condition (2.7).
@@ 2.2. Gauge choice
We fix $\underline{u}$ on the initial outgoing cone $C_{u_{0}}$ by requiring that $\underline{u}=2r+u_{0}$ on $C_{u_{0}}$, and it suggests
@@
As a consequence, in view of $r=0$ on $\Gamma$, it follows that
@@ 2.3. More equations
$\partial_{\underline{u}}\mu={}$ $-\Omega^{2}h\frac{\mu}{r}+\Omega^{2}h\frac{\Phi^{2}}{2r}+\left(1-\mu\right)\Omega^{2}h\frac{\alpha^{2}}{r},$ (2.14a)
$\partial_{u}\mu={}$ $-\underline{h}\frac{\mu}{r}+\underline{h}\frac{\Phi^{2}}{2r}-h\frac{(\partial_{u}w)^{2}}{r}.$ (2.14b)
@@
3. The $H^{1}$ case
We aim to establish an extension theorem for the solution to the SSEYM system in the class of $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$. Recall (1.5) for the definition of the $H^{1}$ norm, where the fixed constant $b$ is given in (1.4).
3.1. Extension theorem in $H^{1}$ space
@@
Given initial data $\alpha|_{C_{u_{0}}}$, assume the initial energy
$\int_{C_{u_{0}}\cap\{u_{0}\leq\underline{u}\leq\underline{u}_{0}\}}\frac{1}{r^{1+b}}\left(\alpha^{2}+(\alpha^{\prime})^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r,\quad\text{for any}\,\,\,\underline{u}_{0}>u_{0}$
is finite. Suppose we have in the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ a solution to the SSEYM system $(r,\,m,\,\alpha)\in C^{1}\times C^{1}\times H^{1}$ (if $b>0$) or $(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}$ (if $b=0$). There is a constant $\varepsilon>0$ such that if
@@
then the following conclusions hold.
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. The solution $(r,\,m,\,\alpha)$ extends as a $C^{1}\times C^{1}\times H^{1}$ solution on $\mathcal{D}(u_{0};\,\underline{u}_{1})$, for some $\underline{u}_{1}>\underline{u}_{\ast}$. Moreover, the solution obeys the estimates $\mathcal{D}(u_{0};\,\underline{u}_{1})$
$|r^{-1}\hat{w}|+|\alpha|+|\partial_{u}w|\lesssim{}$ $r^{b},$ (3.1)
$|r^{-1}\hat{w}|+|\alpha|\lesssim{}$ $\varepsilon^{2},$ (3.2)
and the energy bound holds
$\int_{C_{u}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{1}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r$
$+\int_{\underline{C}_{\underline{u}}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u<C.$ (3.3)
Case II: $b=0$. There exists $\underline{u}_{1}>\underline{u}_{\ast}$ such that the solution extends to $\mathcal{D}(u_{0};\,\underline{u}_{1})$ in the class of $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$, and the estimates (3.2)–(3.1) hold with $b=0$.
@@
Let us introduce some conventions for this section. For any $u_{0}<u_{1}\leq\underline{u}_{\ast}$, we recall the definition of the domain $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ (1.2) that is bounded by $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$. We make the following simplification
@@
and denote the truncated null cones by
@@
Besides, we abbreviate the outgoing cone emanating from the vertex $C_{u}^{[u,\,\underline{u}]}$ by $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ by $\underline{C}_{\underline{u}}$, whenever the interval of integration is clear in the context.
For the fixed constant $b$ (1.4), we define the energy norms,
$E_{0}(u,\,\underline{u}):={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(r^{-2}\hat{w}^{2}+\alpha^{2}\right)\mathrm{d}\underline{u}^{\prime},$
$E_{1}(u,\,\underline{u}):={}$ $E_{0}(u,\,\underline{u})+\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}(\alpha^{\prime})^{2}\mathrm{d}\underline{u}^{\prime},$
$\underline{E}_{0}(u,\,\underline{u}):={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u^{\prime},$
$\underline{E}_{0}^{\prime}(u,\,\underline{u}):={}$ $\underline{E}_{0}(u,\,\underline{u})+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\alpha^{2}\mathrm{d}u^{\prime}.$ (3.5)
@@
$E_{0}={}$ $\int_{0}^{r}\frac{r^{-2}\hat{w}^{2}+\alpha^{2}}{r^{1+b}}\mathrm{d}r|_{C_{u}},$
$E_{1}={}$ $\int_{0}^{r}\frac{(\alpha^{\prime})^{2}+\alpha^{2}+r^{-2}\hat{w}^{2}}{r^{1+b}}\mathrm{d}r|_{C_{u}},$
$\underline{E}_{0}={}$ $\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
$\underline{E}_{0}^{\prime}={}$ $\underline{E}_{0}+\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
@@
3.1.1. Energy estimates theorem
@@
the following statements are concluded.
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. The solution $(r,m,\alpha)\in C^{1}\times C^{1}\times H^{1}$ in $\mathcal{D}$ and it holds that
@@
Case II: $b=0$. The solution $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$ in $\mathcal{D}$ and both of (3.6) and (3.8) hold with $b=0$.
Remark 3.3.
The energy arguments leading to Lemma 3.2 are closed in $H^{1}$ space. This fact enables us to prove a local well-posedness theorem with $H^{1}$ data, see Appendix A.
3.1.2. Preliminary estimates
@@
The fact $\underline{h}<0$ follows from the monotonicity of $\Omega^{-2}\underline{h}$ in $u$ (see (2.4b)) and the initial requirement $\Omega^{-2}\underline{h}|_{u_{0}}<0$, and moreover, $h>0$ is due to the smallness of $\mu$ which indicates $01-\mu=-\underline{h}h>$ 0.
3.1.3. Continuity argument
We make the following bootstrap assumption. Let $\Lambda$ be a large constant that will be determined (depending on the initial energy). We assume
$|r^{-1}\hat{w}|+|\alpha|\leq\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.10)
We will show that the same inequality with $\Lambda$ being replaced by $\frac{\Lambda}{2}$ holds.
@@ 3.1.3. Continuity argument
$|\Phi|\lesssim\varepsilon\Lambda,\,\,|w^{2}-1|\lesssim\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.11)
3.1.4. Main energy estimates
@@
Let us start with the first order energy identities. We multiply $2\partial_{\underline{u}}w$ on the Yang–Mills equation (2.5) to obtain the identity
@@
We further multiply $\Omega^{-2}h^{-1}r^{-1-b}$ on (3.12), then
$\partial_{u}\left(r^{-1-b}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}\right)-\partial_{u}r^{-1-b}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}$
$\qquad-\partial_{u}(\Omega^{-2}h^{-1})\frac{(\partial_{\underline{u}}w)^{2}}{r^{1+b}}+\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3+b}}Q^{2}\right)-\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3+b}}\right)Q^{2}=0.$
@@
$\int_{C_{u}}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}-(1+b)\frac{\alpha^{2}}{r^{2+b}}+b\mu\frac{\alpha^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Note that, among the spacetime integrals on the left hand side of (3.1.4), there is a negative term, which prevents us from deducing any energy bound at this stage.
@@
Multiplied with $hr^{-1-b}$, (3.14) is turned into the following identity with the assistance of (2.4a) and (2.4d),
$\partial_{\underline{u}}\left(h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)+(1+b)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+2\Omega^{-2}(\partial_{\underline{u}}w)^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}$
$\quad+\partial_{u}\left(\frac{\Omega^{2}h}{2}\frac{Q^{2}}{r^{3+b}}\right)+\frac{3+b}{2}\Omega^{2}\underline{h}h\frac{Q^{2}}{r^{4+b}}+\frac{\Omega^{2}}{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{Q^{2}}{r^{3+b}}=0.$
@@
$\int_{\underline{C}_{\underline{u}}}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\frac{\Omega^{2}h}{2}\frac{\Phi^{2}}{r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left((1+b)+2\alpha^{2}\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(-\frac{3+b}{2}+\left(2+\frac{b}{2}\right)\mu\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Analogous to (3.1.4), the above identity has a nonpositive spacetime integral on the left hand side and therefore fails to conclude any energy bound. What is more, cooperating the energy identity (3.1.4) with (3.1.4) does not give any bound for the first order energy neither.
In what follows, we proceed to investigate the second order energy identities, which we expect to cancel those terms with unfavourable signs in (3.1.4) and (3.1.4).
By virtue of (2.3), there is
$\partial_{u}\alpha^{\prime}={}$ $\frac{\underline{h}}{r}\alpha^{\prime}-2w^{2}\frac{h^{-1}}{r}\alpha+2w\frac{h^{-1}}{r}\Phi$
@@
where we denote the linear terms on the first line as follows
@@
Before going into the lengthy calculations for the second order energy estimates, here we sketch the main idea of proof. To obtain an energy identity from (3.1.4), we will multiply on this equation
$r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}=r^{-b}\partial_{\underline{u}}\alpha$
@@
and integrate over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$. Regarding to the first terms on the right hand side of the first line of (3.1.4),

### v2

@@ 1.3. Outline of this paper
This paper is organized as follows. In section 2, we formulate the SSEYM system. Section 3 is devoted to establishing the extension theorem and Cauchy stability in $H^{1}$ space. The proof of local existence theorem is presented in Section 4, while an extension principle away from the axis is provided in Appendix A.
@@ 1.4. Acknowledgement
J.L. is supported by National Key R&D Program of China (No. 2022YFA1005400) and NSFC (No. 12326602, 12141106). J.W. is supported by NSFC (No. 12271450).
@@ 2.1. Spherically symmetric Einstein Yang–Mills equations
In this section, we formulate the SSEYM system with specific gauge assumption.
On a spherical symmetric spacetime $(\mathcal{M},g)$, there exists an isometric action of the rotation group $SO(3)$, whose orbits are two-dimensional spheres. Define the axis $\Gamma$ of $(\mathcal{M},g)$ as the set of fixed points of the $SO(3)$ action. We introduce the double null coordinates $(u,\underline{u})$, and denote the level sets by $C_{u}$ and $\underline{C}_{\underline{u}}$, representing the incoming and outgoing null cones respectively. The intersection $S_{\underline{u},u}=C_{u}\cap\underline{C}_{\underline{u}}$ is a two-dimensional sphere. The area function $r$ is defined by
@@
The metric on $\mathcal{M}$ takes the form
@@
where $\mathrm{d}\sigma_{S^{2}}$ is the standard metric on $S^{2}$ and the lapse function $\Omega$ satisfies $2\Omega^{2}=g(\partial_{u},\partial_{\underline{u}})$.
@@
with $\tau_{i},\,\tau_{j}]=\epsilon_{ijk}\tau_{k}$, $\epsilon_{ijk}$ being the Levi-Civita symbol, $\{i,j,k\}\subset\{1,2,3\}$. We assume the spherically symmetric purely magnetic ansatz [4, 6] for the Yang–Mills potential,
@@
$={}$ $\mathrm{d}w\wedge\left(\tau_{1}\mathrm{d}\theta+\sin\theta\tau_{2}\mathrm{d}\phi\right)-\left(1-w^{2}\right)\tau_{3}\sin\theta\mathrm{d}\theta\wedge\mathrm{d}\phi.$
As followed from the ansatz (2.2), $F_{u\underline{u}}=0$, indicating that the electric charge vanishes; The angular components are given by
$F_{AB}=\frac{w^{2}-1}{r^{2}}\tau_{3}\varepsilon_{AB},$
where capital Latin letters $\{A,B\}$ refer to indices on the 2-sphere $S^{2}$. We define the magnetic charge within the sphere $S_{\underline{u},u}$ as
@@
The vacua of the Yang–Mills field correspond to the values $w=\pm 1$.
The Einstein Yang–Mills system takes the form
@@
where $\hat{D}_{\mu}=D_{\mu}+[A_{\mu},\cdot\,]$ denotes the gauge covariant derivative, $\cdot$$\cdot\,$] is the Lie bracket on $su(2)$, and the energy momentum tensor of Yang–Mills field is given by
@@
We now present the spherical symmetric Einstein Yang–Mills system (SSEYM). With the ansatz (2.2) and the double gauge, the SSEYM system reduces to the following equations for the variables $(r,\Omega,w)$,
@@
A characteristic initial data set for the SSEYM system consists of a triplet $(r,\Omega,\partial_{r}w)|_{C_{u_{0}}}$ prescribed on an outgoing cone $C_{u_{0}}$, as illustrated in Remark 1.1. In addition to the initial condition, the SSEYM system is incorporated with the following boundary condition on the axis $\Gamma$
@@
The condition $Q|_{\Gamma}=0$ is essential to ensure the regularity of the Yang–Mills equation (2.5) on the axis.
We can reformulate the SSEYM system by eliminating $\Omega$ and introducing the mass function $m$. The resulting equivalent system for the variables $(r,m,w)$ comprises the equations
@@
together with the boundary condition (2.7).
@@ 2.2. Gauge choice
We fix the coordinate $\underline{u}$ on the initial outgoing cone $C_{u_{0}}$ by requiring that $\underline{u}=2r+u_{0}$ on $C_{u_{0}}$, which suggests the gauge condition
@@
Since $r=0$ on $\Gamma$, it follows that
@@ 2.3. More equations
$\partial_{\underline{u}}\mu={}$ $-\Omega^{2}h\frac{\mu}{r}+\Omega^{2}h\frac{\Phi^{2}}{r}+2\left(1-\mu\right)\Omega^{2}h\frac{\alpha^{2}}{r},$ (2.14a)
$\partial_{u}\mu={}$ $-\underline{h}\frac{\mu}{r}+\underline{h}\frac{\Phi^{2}}{r}-2h\frac{(\partial_{u}w)^{2}}{r}.$ (2.14b)
@@
3. Extension theorem in $H^{1}$ space
We aim to establish an extension theorem for solutions to the SSEYM system in the class of $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$. Recall (1.4) for the definition of $H^{1}$ norm.
@@ 3. Extension theorem in $H^{1}$ space
Given initial data $\alpha|_{C_{u_{0}}}$, with finite initial energy
$\int_{C_{u_{0}}\cap\{u_{0}\leq\underline{u}\leq\underline{u}_{0}\}}\frac{1}{r}\left(\alpha^{2}+(\alpha^{\prime})^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r,\quad\text{for any}\,\,\,\underline{u}_{0}>u_{0},$
suppose we have on the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ a solution to the SSEYM system $(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}$. There is a constant $\varepsilon>0$ such that if
@@
then the solution extends to $\mathcal{D}(u_{0};\,\underline{u}_{1})$, for some $\underline{u}_{1}>\underline{u}_{\ast}$, with the same regularity $C^{1}\times C^{0}\times H^{1}$.
Moreover, the solution obeys the estimates on $\mathcal{D}(u_{0};\,\underline{u}_{1})$
$|r^{-1}\hat{w}|^{2}\lesssim\varepsilon^{2},\quad\text{and}\quad|\alpha|^{2}\lesssim\varepsilon,$ (3.1)
and the energies remain bounded
$\int_{C_{u}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{1}{r}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r$
$+\int_{\underline{C}_{\underline{u}}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{h^{-1}}{r}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u<C.$ (3.2)
@@
3.1. Conventions
Let us introduce some conventions for this section. For any $u_{0}<u_{1}\leq\underline{u}_{\ast}$, we recall the definition of the domain $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ from (1.2), which is bounded by $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$. We make the following simplification
@@ 3.1. Conventions
and denote the truncated null cones as
@@
Furthermore, we abbreviate the outgoing cone segment emanating from the vertex $C_{u}^{[u,\,\underline{u}]}$ as $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ as $\underline{C}_{\underline{u}}$, when the interval of integration is clear in the context.
We define the energy norms,
$E_{0}(u,\,\underline{u}):={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r}\left(r^{-2}\hat{w}^{2}+\alpha^{2}\right)\mathrm{d}\underline{u}^{\prime},$
$E_{1}(u,\,\underline{u}):={}$ $E_{0}(u,\,\underline{u})+\int_{C_{u}}\frac{\Omega^{2}h}{r}(\alpha^{\prime})^{2}\mathrm{d}\underline{u}^{\prime},$
$\underline{E}_{0}(u,\,\underline{u}):={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u^{\prime},$
$\underline{E}_{0}^{\prime}(u,\,\underline{u}):={}$ $\underline{E}_{0}(u,\,\underline{u})+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\alpha^{2}\mathrm{d}u^{\prime}.$ (3.4)
@@
$E_{0}={}$ $\int_{0}^{r}\frac{r^{-2}\hat{w}^{2}+\alpha^{2}}{r}\mathrm{d}r|_{C_{u}},$
$E_{1}={}$ $\int_{0}^{r}\frac{(\alpha^{\prime})^{2}+\alpha^{2}+r^{-2}\hat{w}^{2}}{r}\mathrm{d}r|_{C_{u}},$
$\underline{E}_{0}={}$ $\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}}{r}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
$\underline{E}_{0}^{\prime}={}$ $\underline{E}_{0}+\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{\alpha^{2}}{r}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
@@
3.2. Energy estimates theorem
@@
the solution $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$ in $\mathcal{D}$. Moreover, it holds that
@@
Remark $3.$
The energy arguments leading to Lemma 3.2 are closed in the $H^{1}$ space. This fact enables us to prove a local well-posedness theorem with $H^{1}$ data, as detailed in Section 4.
3.3. Preliminary estimates
@@
The fact $\underline{h}<0$ follows from the monotonicity of $\Omega^{-2}\underline{h}$ in $u$ (see (2.4b)) and the initial requirement $\Omega^{-2}\underline{h}|_{u_{0}}<0$, and moreover, $h>0$ is due to the smallness of $\mu$ which indicates $1-\mu=-\underline{h}h>0$.
3.4. Continuity argument
We make the following bootstrap assumption. Let $\Lambda$ be a large constant that will be determined (depending on the initial energy) later. We assume
$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}\leq\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.8)
and will show that the same inequality with $\Lambda$ being replaced by $\frac{\Lambda}{2}$ holds.
@@ 3.4. Continuity argument
$|\Phi|^{2}\lesssim\varepsilon\Lambda,\,\,\text{and}\,\,|w^{2}-1|^{2}\lesssim\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.9)
3.5. Main energy estimates
@@
Let us begin with the first order energy identities. We multiply $2\partial_{\underline{u}}w$ on the Yang–Mills equation (2.5) to obtain the identity
@@
We further multiply $\Omega^{-2}h^{-1}r^{-1}$ on (3.10), then
$\partial_{u}\left(r^{-1}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}\right)-\partial_{u}r^{-1}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}$
$\qquad-\partial_{u}(\Omega^{-2}h^{-1})\frac{(\partial_{\underline{u}}w)^{2}}{r}+\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3}}Q^{2}\right)-\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3}}\right)Q^{2}=0.$
@@
$\int_{C_{u}}\Omega^{2}h\frac{\alpha^{2}}{r}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}\frac{\Phi^{2}}{2r}\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(\frac{3}{2}\frac{\Phi^{2}}{r^{2}}-\frac{\alpha^{2}}{r^{2}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$={}$ $\int_{C_{u_{0}}}\Omega^{2}h\frac{\alpha^{2}}{r}\mathrm{d}\underline{u}.$ (3.11)
Note that, among the spacetime integrals on the left-hand side of (3.5), there appears a negative term, which obstructs the derivation of energy bounds at this stage.
@@
Multiplying by $hr^{-1}$ and using equations (2.4a), (2.4d), we transform (3.12) into the following identity,
$\partial_{\underline{u}}\left(h\frac{(\partial_{u}w)^{2}}{r}\right)+\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2}}+2\Omega^{-2}(\partial_{\underline{u}}w)^{2}\frac{(\partial_{u}w)^{2}}{r^{2}}$
$\quad+\partial_{u}\left(\frac{\Omega^{2}h}{2}\frac{Q^{2}}{r^{3}}\right)+\frac{3}{2}\Omega^{2}\underline{h}h\frac{Q^{2}}{r^{4}}+\frac{\Omega^{2}}{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{Q^{2}}{r^{3}}=0.$
@@
$\int_{\underline{C}_{\underline{u}}}h\frac{(\partial_{u}w)^{2}}{r}\mathrm{d}u+\int_{C_{u}}\Omega^{2}h\frac{\Phi^{2}}{2r}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(1+2\alpha^{2}\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(-\frac{3}{2}+2\mu\right)\Omega^{2}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$
$={}$ $\int_{C_{u_{0}}}\Omega^{2}h\frac{\Phi^{2}}{2r}\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\Omega^{2}\Phi^{2}\frac{\Phi^{2}}{2r^{2}}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.13)
Analogous to (3.5), the identity above contains a nonpositive spacetime integral on the left-hand side, which prevents it from yielding an energy bound. What is more, cooperating the energy identity (3.5) with (3.5) does not give any bound for the first order energies neither.
To address this, we now derive the second order energy identities, which are designed to absorb the terms with unfavorable signs in (3.5) and (3.5).
By virtue of (2.3), we have
$\partial_{u}\alpha^{\prime}={}$ $\frac{\underline{h}}{r}\alpha^{\prime}-2w^{2}\frac{h^{-1}}{r}\alpha+2w\frac{h^{-1}}{r}\frac{Q}{r}$
@@
where we mark the linear terms on the first line as below
@@
Before proceeding to the detailed second order energy estimates, we outline the main idea of the proof. To obtain an energy identity from (3.5), we multiply the equation by
$r^{-1}\cdot\Omega^{2}h\alpha^{\prime}=\partial_{\underline{u}}\alpha$
@@
and integrate over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$. Regarding to the first terms on the right-hand side of the first line,